OpenStudy (anonymous):

identify the outlier in the data then find the MEAN when the outlier is included and when it's not included 68,63,66,65,87,69,61,86,82,28. @phi

OpenStudy (anonymous):

@undeadknight26 @pooja195 @popcornfun @babyboy2332 @Kitten_is_back @imqwerty @Angelofdarkness80 @amitrashok @Alphabet_Sam @Anaise @adelera @Agl202

OpenStudy (anonymous):

will medal

undeadknight26 (undeadknight26):

The outlier is the number that is very very far from the others.

OpenStudy (anonymous):

so 87?

undeadknight26 (undeadknight26):

86 is close to 87 though! So they are not far from each other.

undeadknight26 (undeadknight26):

Try putting them in order from least to greatest for me :)

OpenStudy (anonymous):

28,61,63,65,66,68,69,82,86

undeadknight26 (undeadknight26):

Can you tell me which number here is the farthest away from all the rest? It's uaully at the ends :)

OpenStudy (anonymous):

86

OpenStudy (anonymous):

i think it's Outlier: 28 Mean with outlier: add the numbers up and divide by the amount of numbers there are... 742/11=67.45455 Mean without outlier: 714/10=71.4

OpenStudy (anonymous):

idk if im right tho

undeadknight26 (undeadknight26):

28,61,63,65,66,68,69,82,86 The outlier would be 28.

OpenStudy (anonymous):

yeah that what i put Outlier: 28 Mean with outlier: add the numbers up and divide by the amount of numbers there are... 742/11=67.45455 Mean without outlier: 714/10=71.4

OpenStudy (anonymous):

am i right

undeadknight26 (undeadknight26):

If you put the outlier as 28 you should be right :)

OpenStudy (anonymous):

thankyou @undeadknight26